OpenStudy (mikee):

A sofa is on sale for $425.00 , which is 32% less than the regular price. What is the regular price?

OpenStudy (anonymous):

$425=X-0.32X $425=0.68X X = $425/0.68 X = $625

OpenStudy (mikee):

Ty Nancy

OpenStudy (anonymous):

u r wel com If I deserve medal then give me one

OpenStudy (mikee):

How do i give you one? You deserve on for helping me with my last problem.

OpenStudy (anonymous):

next to my name , click on give medal

OpenStudy (mikee):

Done and Ty again :)

OpenStudy (anonymous):

u did ,t y for medal

OpenStudy (mikee):

Yw Nancy :)

OpenStudy (anonymous):

u we com